# Two Sides of the Same Coin: Large-scale Measurements of Builder and Rollup after EIP-4844

### v2

@@ Appendix C Greedy Algorithm
2for $tx_{j}$ in $\mathbb{T}_{r}$ do
3       $\text{ratio}\leftarrow\text{$v_{j}$}/\text{$s_{j}$}$;
@@
[…]
11 for $tx_{j}$ in $\mathbb{T}_{r}$ do
12       if $\mathrm{total\_size+\text{$s_{j}$}}\leq\beta\times 128$ then
13             $\mathrm{total\_size}\leftarrow\mathrm{total\_size}+\text{$s_{j}$}$;
14             $\mathrm{total\_tips}\leftarrow\mathrm{total\_tips}+\text{$v_{j}$}$;
15             $\mathrm{selected\_list}.\text{append}(\text{$tx_{j}$})$;
@@ 14             $\mathrm{total\_tips}\leftarrow\mathrm{total\_tips}+\text{$v_{j}$}$;
[…]
@@
We use a greedy algorithm as shown in Algorithm 1, considering the size and priority fee per transaction, calculating the priority fee ratio for each transaction and sorting by the ratio before processing.
Appendix D Invisible Profits for Builders
@@
The profit of type-3 transactions size in the public mempool among builders.
Figure 12. The profit of type-3 transactions size in the public mempool among builders. Potential profits that can be extracted from the public mempool exceed those from type-3 transactions.
In Figure 12, we plot the results for six months in different colors among the main builders. The right y-axis of Figure 12 shows the profits that builders can extract from the public mempool, corresponding to the blue bars, while the left y-axis represents the transactions that are actually extracted from type-3 transactions, corresponding to the green bars. It reveals that builder can extract more profit in public mempool, but instead they chose to include the much cheaper type-3 transaction into block. The cumulative profits of other transactions for each builder are overwhelming compared to type-3 transactions, which explains that type-3 transactions give significantly lower priority fees than the other transactions, and do not compensate at all for the impact of their size. In addition, we observe that rsync builder may choose fewer type-3 transactions in their blocks because Vanilla builders have a lower market share but are more profitable from both type-3 transactions and other transactions from public mempool.
@@
Figure 13. The relationship between delays and maxPriorityFeePerGas among rollups. Similar price from rollups have uneven inclusion latency.
Figure 13 illustrates the average delay and average priority fee of different rollups in August. The y- and x-axes indicate the average latency and priority fee, respectively, while each point denotes a particular rollup. In the context of the builder sorting algorithm being absolutely efficient, the inclusion delay of type-3 transactions with the same priority fee should be close. However, in Figure 13, rollups that offer close to average priority fees have completely different inclusion delays. For example, Optimism has a higher priority fee than Zircuit, Kroma, and Arbitrum, but the latter still has a shorter waiting time for type-3 transactions than Optimism. On top of that, Kroma and Arbitrum, etc. have transaction latencies that are very comparable to Rari, but their transaction priority fees are just half that of Rari.
Appendix E Private Channels between Builder and Rollups.
We compare on-chain data with data recorded by public mempool and label data that bypassed public mempool as private transactions from EIP-4844 to August. Our final private type-3 transaction dataset consists of 3,153 private transactions, of which 826 transactions come from a rollup called Lisk, which accounts for 98% of the type-3 transactions issued by this roll-up. Interestingly, 1,405 of these 1,484 private transactions from Taiko were built into the block by the Titan builder to be on the main net, and 795 transactions from Lisk also built into Titan’s blocks.
Table 3. The Top 5 private type-3 transaction senders
Address Rollup Counts
0x000000……663bcbe0 Taiko 1484
0xa6Ea2f……9cd6fe24 Lisk 826
0x2C169D……fe260ca7 StarkNet 179
0x0D3250……f790ec99 zkSync 128
0xcF2898……0b4cfc8f Scroll 147
Appendix F Reduction in Public Mempool
@@
Figure 14. The inclusion delay for type-3 and non type-3 transactions over time.
A smaller public transaction block size may lead to an increased transaction inclusion delay. To evaluate it, we display the overall transaction latency in Figure 14. Nevertheless, transactions in the public mempool did not experience extended inclusion delays, and the overall trend consistently decreased. The decrease in public mempool transactions explains why there hasn’t been an increase in latency. In fact, there were 28,012,234 transactions in the public mempool in August, while in March, there were 41,500,537 transactions, highlighting a decrease of 13,488,303 transactions.
Table 4. Granger causality test of block size and mempool size
$lag_{x}=lag_{y}$ F-value p-value
1 35011.6924 0.00
2 26616.7382 0.00
3 20894.1280 0.00
4 17072.5844 0.00
5 14266.8419 0.00
We use Pearson’s correlation coefficient and the Granger causality test to further evidence the correlation between the size builder extracts from the public mempool and the quantity of the public mempool. The Pearson correlation coefficient is 0.4335 with a p-value of 0.00. The p-value indicates the relationship between the overall size of transactions extracted by the builder from the mempool and the decreasing quantity of mempool. We also perform a Granger causality test in Table 4, indicating that a decrease in size correlates with a reduction in the number of transactions in the public mempool. The reduction in the public mempool could be associated with EIP-4844. This is because more transaction senders are choosing to use the lower-priced Layer-2 after EIP-4844 instead of the Ethereum mainnet (Pokorny, 2024).

### v3

@@ 4.3. Efficiency of Builders
2for $m_{i}$ in $\mathbb{T}_{r}$ do
3  $\text{ratio}\leftarrow\text{$v_{i}$}/\text{$s_{i}$}$;
@@
9 $\text{total\_tips}\leftarrow 0$;
10 $\text{selected\_list}\leftarrow[]$;
11 for $m_{i}$ in $\mathbb{T}_{r}$ do
12  if $\mathrm{total\_size+\text{$s_{i}$}}\leq q\times 128$ then
13     $\mathrm{total\_size}\leftarrow\mathrm{total\_size}+\text{$s_{i}$}$;
14     $\mathrm{total\_tips}\leftarrow\mathrm{total\_tips}+\text{$v_{i}$}$;
15     $\mathrm{selected\_list}.\text{append}(\text{$m_{i}$})$;
@@ 14     $\mathrm{total\_tips}\leftarrow\mathrm{total\_tips}+\text{$v_{i}$}$;
21 return $\text{total\_tips},\text{selected\_list}$;
@@
Table 1. The overall inefficient block distribution over time.
March April May June July August Overall
Inefficient Block 26,450 41,779 47,351 31,540 27,965 23,139 198,224
Total Block 47,604 84,631 105,382 139,529 145,726 149,515 672,388
Inefficient Proportion 55.56% 49.37% 44.93% 22.60% 19.19% 15.48% 29.48%
To measure whether builders’ new strategies are efficient or not, we employ a greedy algorithm considering size and profit to measure whether the priority fee given to each blob can match the priority fee that can be extracted in the public mempool during the concurrent period. Note that before being executed, only the maximum priority fee, i.e. maxPriorityFeePerGas field, can be seen. This represents the profit for builders in one transaction. We use this value to calculate the potential profit for builders. To filter out transactions that builders have simulated as likely to fail, we filter out transactions in the public mempool with a delay of more than 100 seconds to make the calculation more accurate. We also provide the pseudocode of the greedy algorithm. The algorithm considers the size and priority fee per transaction, calculating the priority fee ratio for each transaction and sorting by the ratio before processing.
3(a) plots the profit that builders actually extract from type-3 transactions versus the profit that can be extracted from public mempool in the same period from March to August. We compute the profit difference between the profit of type-3 transactions and the profit that can be extracted from the public mempool, and indicate this value with red and yellow bars. The red bar indicates that this difference is negative, which suggests that the builders could have made more profit from the public mempool, but they chose to include the less profitable type-3 transactions in the block. Conversely, if the color is yellow, it indicates that the value is positive and the builder made more profit from type-3 transactions. It is clear that after EIP-4844, the builder misses significant value of the public mempool. As 3(a) shows, in March, builders can potentially make more than five times more profit from the public mempool as they can from type-3 transactions. However, this loss diminishes over time and reverses in July, when the builder makes more profit from type-3 transactions than it is able to make from the public mempool.
Table 1 illustrates the details from March to August, including the number of efficient and inefficient blocks and their proportion out of the 636,981 blocks containing type-3 transactions. We find that 196,090 blocks are inefficient, which is approximately 29.48%. Encouragingly, the percentage of inefficient blocks dropped from 55.56% immediately after EIP-4844 to 15.48% by August, indicating a trend toward a more efficient block building market. However, 15.48% remains a non-negligible share, underscoring that there is still considerable room for improvement.  3(b) shows the ratio of efficient blocks among major builders during the same period. The y-axis shows the proportion of efficient and inefficient blocks. Beaverbuild has a market share of almost 50%, and is the most efficient compared to other builders. In contrast, non-specialist builders, vanilla builders who do not specialize in block construction, have nearly 30% more inefficient blocks than beaverbuild, and more than half of the blocks they construct are inefficient.
In 4(a), we plot the results for six months in different colors among the main builders. The right y-axis of 4(a) shows the profits that builders can extract from the public mempool, corresponding to the blue bars, while the left y-axis represents the transactions that are actually extracted from type-3 transactions, corresponding to the green bars. It reveals that the builder can extract more profit in public mempool, but instead they chose to include the much cheaper type-3 transaction into the block. The cumulative profits of other transactions for each builder are overwhelming compared to type-3 transactions. This is because type-3 transactions give significantly lower priority fees than the other transactions, and do not compensate for the impact of their size. In addition, we observe that rsync builder may choose fewer type-3 transactions in their blocks and Vanilla builders have a lower market share but are more profitable from both type-3 transactions and other transactions from public mempool.
Answer to RQ1 EIP-4844 significantly altered block construction strategies: the overall block size expanded dramatically by approximately 2.5 times, growing from less than 150 KB to close to 400 KB. Concurrently, the available capacity for non type-3 transactions was reduced roughly fourfold from 300 KB to 80 KB, indicating that type-3 transactions crowded out other transactions. From March to August, 29.48% of the 636,981 blocks containing type-3 transactions were inefficient, with builders missing potential profits from the public mempool. In March alone, builders could have earned over 5× more from the mempool than from type-3 transactions. However, efficiency improved over time, with the share of inefficient blocks dropping from 55.56% post-EIP-4844 to 15.48% by August, driven largely by specialist builders like Beaverbuild, whose inefficiency rate is over 30% lower than that of non-specialist builders.
5. Understanding Rollup Strategies
After the overview in builder market, as outlined in RQ2 in Section 1, we now present a measurement of rollups to obtain a comprehensive understanding of their current strategies toward blob and type-3 transaction usage, as well as their efficiency. In this section, we examine the opposite side of the market, rollup’s strategy. We start by analyzing the different rollup behaviors after EIP-4844. We conduct an analysis about the efficiency of rollups and identify two inefficient practices and measure their impact.
5.1. Type-3 Transactions Usage
After the upgrade, in total 1,336,822 type-3 transactions have been sent to the Ethereum mainnet. 4(b) visualizes the daily share of type-3 transactions by the dominant rollups and a specific role after EIP-4844, in terms of the total number of blob-carrying transactions. We notice that the number of transactions for each rollup varies over time, with some rollups showing significant spikes or drops. One of the most noticeable spikes is in Blobscriptions, which began hitting the market at the end of March, reached its peak, and then began to decrease dramatically. The purpose of Blobscriptions is not to validate rollup data, but rather to publish it. Other rollups show varying increases and decreases in usage, which is related to the choices made by each rollup application. Obviously, the use of type-3 transactions has been increasing after EIP-4844, with more rollup applications using blobs to validate data. Instead of immediately using a blob to transfer data, some rollups (e.g., Taiko) only started adopting type-3 transactions in May of this year. However, there are still many rollup applications (such as Arbitrum and Base) that have been using type-3 transactions since EIP-4844.
(a) The profit of type-3 transactions size in the public mempool among builders.
@@ 5.1. Type-3 Transactions Usage
(b) The landscape of type-3 transactions.
@@
Figure 4. The landscape of type-3 transactions and blob usage.
In  5(a), we plot the blob packing strategy for each mainstream rollup from 13 March to 31 August. The different colors represent the number of blobs carried in type-3 transactions and the Y-axis represents the total number of transactions. It is clear that rollup makes different choices in selecting the number of blobs carried in a type-3 transaction. Of the 343,468 transactions issued by Taiko, it puts only one blob into the transaction, and the same strategy is used by Scroll, Starknet, and Zircuit. We further investigate the reasons for this in Section 6.2, and find that the latency of the transactions that carry only one blob is the lowest. Therefore, it is likely that this strategy is optimizing the latency of the rollup application so that the user’s transactions can be verified the fastest. In contrast, Base and Op Mainnet, which have been using blobs to send data, put as many blobs as possible into a transaction, and most of their transactions carry five to six blobs in a single transaction. In addition, there are some rollups that do not limit how many blobs they carry in a single transaction (such as Linea and Blast).
We also show in 5(b), the percentage of type-3 transactions issued by different rollups that carry the same number of blobs. 69.85 % of the transactions in the network carry only one blob, while a staggering 1,239 transactions carry four blobs. Rollups that send one blob per type-3 transaction dominate. Most rollups use relatively fixed strategies. In addition to that, 2-4 blobs per transaction is a minimal amount.
(a) The blob packing strategies among rollups.
@@
(b) The landscape of blob usage strategies.
[…]
Block size.
Figure 5. On left is the blob number distribution bar chart of type-3 transactions, on right is the relationship between delays and tips among different rollups.
5.2. Inefficient Strategies
(a) The behaviors and amount of inefficient strategies.
[…]
Block size.
(b) The cumulative loss of inefficient strategies.
[…]
Type-3 excluded block size.
Figure 6. On left is the overview of inefficient strategies from rollups. On right is the cumulative loss of Continuous Sending for rollups.
To analyze whether rollups are using type-3 transactions and blobs efficiently, we analyze their blob-carrying strategies and identify two existing inefficient strategies, Continuous Sending and Double Sending, and another usage, blobscription. 6(a) presents a diagram of these inefficient strategies. A Continuous Sending pattern is when the type-3 transaction sender sends more than one transaction in succession that are ultimately contained within the same block. These blobs can be put into a single type-3 transaction instead of multiple ones, which would require duplicate payments. Double Sending occurs when a rollup transmits the identical blob multiple times. A type-3 transaction that carries a second identical blob is invalid because the first successful transaction completes the validation of the blob data. These two strategies account for 7% and 10% of the total transactions, respectively. Additionally, it is important to mention that blobscription is observed in 106,051 transactions in the network, representing 8% of the total. This section evaluates the impact and economic loss to the network of the inefficient strategies, showing the loss of 186.92 ETH and an average delay in network inclusion of 19 seconds.
Economic Cost.  We calculate the transaction fee (including base fee and priority fee) for duplicate transactions in Continuous Sending and Double Sending.  6(b) plots the cumulative losses from the Double Sending and Continuous Sending strategies for rollups after EIP-4844. The left y-axis shows the cumulative loss for these two strategies, and the right y-axis shows the total number of type-3 transactions. We use light blue bars to show the most rollups per month using these two strategies, and the yellow line represents cumulative losses. Since EIP-4844, rollups have lost 186.92 ETH. while Scroll sends the largest volume of transactions using the two inefficient strategies, it is worth noting that the number of these strategies has been decreasing since May, with less than 1 transaction in August. Another finding is that the use of the continuous sending strategy almost stopped in April, this is because Linea sent significantly fewer type-3 transactions in April.
Network Delay.  The presence of a large number of blobscription transactions violates the original intent of the EIP-4844 design and interferes with the normal rollup validation process. Since a block can only carry a maximum of 6 blobs, the massive number of blobscription transactions in the public mempool may cause a delay in the normal type-3 transactions. We further measure the impact of blobscription on network latency. We analyze 100,504 blobscription transaction during a period of burst of blobscription usage. During this period, there were 151,688 type-3 transaction sent from the rollups, and 21,282 of them were affected by blobscription. The results show that the average latency of the affected type-3 transaction was 43 seconds, while the other type-3 transactions only had to wait 24 seconds. The presence of blobscription in the network introduces a delay of nearly 19 seconds to the rollups. Our finding exposes that if a large number of blobs that are not used for data validation are consistently present within the network, this causes a huge impact on the process of uploading type-3 transactions for rollups, which in turn affects the efficiency of the rollup application.
Answer to RQ2 69.85% of type-3 transactions carried only one blob, with rollups like Taiko and Scroll optimizing for latency, while others like Base and Op Mainnet frequently packed 5–6 blobs per transaction for higher data throughput. Our analysis identifies two inefficient blob-carrying strategies — Continuous Sending and Double Sending—used by rollups, both leading to redundant data submissions and unnecessary costs. Combined with 106,051 blobscription transactions, these inefficiencies caused a total loss of 186.92 ETH and an average network inclusion delay of 19 seconds.
6. Understanding Optimal Strategies
The previous sections have separately examined the trading strategies of builders and rollups in the post-EIP-4844 period. The results indicate that block construction strategies adopted by builders become inefficient with the emergence of type-3 transactions, and that the submission strategies of type-3 transactions through rollups also exhibit inefficiencies. To address this, the interaction between builders and rollups is modeled in this section to determine the optimal pricing of type-3 transactions. The resulting analysis provides both the optimal price that builders should select and the preferred blob-carrying strategy that rollups should adopt.
6.1. Optimal Price for Builders
Table 2. Frequently used notations.
Notation Description
$k$ the number of type-3 transactions in one slot
$q$ the number of blobs carried in one type-3 transaction
$\mathbb{N}$ a set of type-3 transaction $\mathbb{N}=\{n_{1},n_{2},\ldots,n_{k}\}$
$P(n)$ the floor price for blockspace usage by type-3 transactions $n$
$R,\tilde{R}$ the revenue of type-3 transaction and its initial revenue without discount.
$\rho(t)$ the time-dependent discount factor of type-3 transaction revenue.
$b(R)$ the pricing strategy of the type-3 transactions with revenue $R$.
$\pi(R)$ the winning probability of the type-3 transaction with revenue $R$
In the game between builders and rollups, builders privately select a subset of transactions to fill the remaining blockspace, and each rollup submits its type-3 transactions without observing the price of others—mirroring the sealed-bid setting. Moreover, the payment rule is effectively first-price: when a transaction is selected, it pays the exact fee it has given. To capture these mechanics, as well as the effect of limited blockspace and the presence of a floor price $P(n)$, which we have shown in Section 4.2, we formalize the selection process as a first-price sealed-bid auction with size constraints and a reserve floor price $P(n)$. For clarity and ease of reference, Table 2 provides a summary of the notations used throughout the paper, including those that have already appeared in earlier sections.
We assume that type-3 transactions $\mathbb{N}$ compete for the remaining space $s^{*}$ in each block. At this stage, type-3 transactions $\mathbb{N}$ are assumed to compete for the remaining blockspace $s^{*}$. As defined in Section 4.2, a type-3 transaction is represented as $n$. The revenue of type-3 transaction is denoted as $R$. It should be noted that type-3 transactions are time sensitive. For instance, fraud proofs in optimistic rollups must be submitted within a strict challenge window, while validity proofs in ZK-rollups directly determine the finalization of batched Layer-2 transactions on the mainnet (Scroll, 2025; Arbitrum, 2025; Taiko, 2025). In practice, rollups often prioritize certain type-3 transactions—such as validity proofs, fraud proofs, and withdrawal finalization proofs—due to their direct impact on security and user experience. Therefore, a time-dependent discount factor $\rho(t)$ is introduced, where $t$ denotes discrete block slots, i.e., each increment of $t$ corresponds to one block proposal opportunity. The discount captures the fact that the economic value of a type-3 transaction decays as it waits to be included in later slots. Under this formulation, the revenue can be expressed as
(3) $R(q,t)=\rho(t)\tilde{R},128q\leq s^{*},$
where $\tilde{R}$ denotes the revenue at the initial time period $t=0$. The function $\rho(t)$ is strictly decreasing, with $\rho(0)=1$ and $\rho^{\prime}(t)<0$.
Under this setting, each type-3 transaction $n$ must strategically determine its price pp—balancing the inclusion probability against utility maximization. If all type-3 transactions adopt a strictly increasing and differentiable pricing strategy, $p=b(R)$, a type-3 transaction with revenue $R$ wins if and only if its price exceeds all other price.
Further, unlike ordinary auction games, the selection probability of type-3 transactions may also depend on transaction size. Hence, before defining the winning probability, it is necessary to evaluate whether block size constraints affect the inclusion probability of type-3 transactions. Figure 7 presents the relationship between the number of blobs (transaction size) and the average inclusion latency of type-3 transactions after EIP-4844, with a fitted trend for the observed delays. The y-axis represents the inclusion delay, and the x-axis indicates the number of blobs per transaction. Bar colors, ranging from blue to green, correspond to transaction counts. Empirical evidence shows that, after June, type-3 transactions containing between 1 and 5 blobs incur negligible additional latency. Under these conditions, it is reasonable to assume that sufficient blockspace is available, and block size constraints do not materially affect the selection probability. Consequently, the probability of inclusion can be simply expressed as $\pi(R)=F(R)^{n-1}$. The analysis first focuses on a single time period setting, assuming that the revenue $R_{t_{0}}=\rho(t_{0})\tilde{R}$ follows a distribution characterized by the cumulative density function $F(\cdot)$. With the reserve price $P(n)$, the type-3 transaction’s expected utility is
(4) $\mathbb{E}\left[U_{t_{0}}\right]=\Pr\left[\mathrm{win}|R\right]\left(R_{t_{0}}-p\right)=\begin{cases}\pi(R_{t_{0}})\left(R_{t_{0}}-b(R_{t_{0}})\right),\quad&b(R_{t_{0}})\geq P(n).\\ 0,\quad&b(R_{t_{0}})<P(n).\end{cases}$
At symmetric equilibrium, each type-3 transaction selects $p=b(R)$ to maximize $U$. Differentiation with respect to $p$, together with the envelope condition, yields the standard first-order condition (Krishna, 2009) as
(5) $p=\rho(t_{0})\left(\tilde{R}-\frac{\int_{0}^{R_{t_{0}}}\pi(x)dx}{\pi(R_{t_{0}})}\right).$
For tractability, $R$ is always assumed to be independently drawn from a uniform distribution in auction games, i.e., $R\sim U(0,R_{max})$ (Krishna, 2009; Qin et al., 2022). We assume that $k$ and $U(0,R_{max})$ are prior knowledge of builders and rollups. Under this simplifying assumption, the price for type-3 transactions accepted by builders in period $t_{0}$ satisfies
(6) $p_{t_{0}}^{*}=\max\left(P(n_{j}),\frac{k-1}{k}\max_{j}\rho(t_{0})\tilde{R}\right),1\leq j\leq k.$
This represents the builder’s optimal selection strategy when the price of type-3 transactions reach the Bayesian Nash equilibrium. It can be observed that, as the number of type-3 transactions increases, the price obtained by the builders increases and gradually approaches the revenue that the rollups derive from these transactions.
6.2. Blob-carrying Strategy for Rollups
While analyzing the optimal price selection strategy of the builder in Section 6.1, the corresponding pricing behavior of rollups has been implicitly addressed. The discussion now shifts to another challenge in submitting type-3 transactions through rollups: the selection of an appropriate blob-carrying strategy. When multiple blobs need to be submitted within a single period, as described in Section 5.1, the rollup must determine how many blobs are bundled into each type-3 transaction. This corresponds to a classical integer partition problem. In practice, rollups such as Taiko, Scroll, and Starknet tend to split these blobs into several type-3 transactions, whereas Arbitrum and Base choose to aggregate all blobs into a single type-3 transaction. A comparison of these strategies is therefore required. We extend the single-period first-price sealed-bid auction model introduced earlier to a multi-round framework.
[…]
Inclusion delay and blob number.
Figure 7. Inclusion delay and blob number.
The analysis starts with a single rollup submitting two blobs, serving as the foundational scenario for further cases. These blobs generate revenues $R_{1}$ and $R_{2}$. We continue our analysis from the previous single time-period game. After failure, transactions enter the next phase and the blob revenue decays to $\rho(k)$, where $k$ is the number of failed rounds. Therefore, the expected utility of multiple periods becomes a geometric distribution form as
(7) $\mathbb{E}[U]=\frac{\pi(\tilde{R)}}{n}\Phi\left(\pi(\tilde{R})\right),\quad\Phi\left(\pi(\tilde{R})\right)=\sum_{k=1}^{\infty}(1-\pi(\tilde{R}))^{k}\rho(k).$
For convenience, we define
(8) $g(R)=\frac{\pi(R)}{n}R\Phi\bigl{(}\pi(R)\bigr{)}.$
With this notation, the expected utility of bundling two blobs into a single type-3 transaction is $\mathbb{E}[U_{\mathrm{bundle}}]=g(R_{1}+R_{2})$, whereas the expected utility of submitting them separately becomes $\mathbb{E}[U_{\mathrm{split}}]=g(R_{1})+g(R_{2})$. The comparison between these strategies thus reduces to examining the curvature of $g(R)$.
From the second derivative of $g(R)$, it is observed that when $\rho(k)$ decreases slowly, $\pi(R)$ grows rapidly, and $g(R)$ exhibits convexity with $\pi^{\prime\prime}(R)$ remaining positive under a uniform distribution. It means that bundling multiple blobs into a single type-3 transaction is generally the more advantageous strategy. Only in cases where $\rho(k)$ decays sharply—indicating high delay sensitivity and reduced curvature of $\pi(R)$$g(R)$ becomes nearly linear or slightly concave, making a split-submission strategy more beneficial. However, empirical evidence in Figure 7 shows that type-3 transactions carrying multiple blobs do not incur significant additional latency. Hence, bundling multiple blobs into a single type-3 transaction is expected to yield higher utility. Starting from the two-blob case analyzed above, the model extends naturally to the general case of multiple blobs, and the conclusion remains valid. From our landscape analysis in Section 5.1, this is exactly the strategy used by Arbitrarum and Base. In conclusion. from our model and derivation, we give the optimal pricing selection strategy of builders and the optimal blob-carrying strategy of rollups, therefore successfully answer the RQ3.
[…]
7. Related Work
Ethereum Upgrades.  Prior works have explored several upgrades related to Ethereum, including the appearance of ERC-20 tokens (Li et al., 2023; Hu et al., 2024b; Gao et al., 2020; Wu et al., 2023; Xia et al., 2021), emergence of NFT (Huang et al., 2024; Ito et al., 2024; Huang et al., 2023), the London hard fork (Liu et al., 2022; Leonardos et al., 2023), and the transition from PoW to PoS (Heimbach et al., 2023b; Silva et al., 2020; Chi et al., 2024; McLaughlin et al., 2023). After these updates, detection frameworks have been proposed to address the security issues posed by cryptocurrency tokens (Hu et al., 2024a; Victor and Weintraud, 2021; Li et al., 2022). Hu et al. (Hu et al., 2023) developed a model to detect Ethereum fraud, while Zhou et al. (Zhou et al., 2024) examined the use of NFT for arbitrage and airdrop acquisition and revealed that it is driven by a small number of dominant players. Regarding the Ethereum merge event, Heimbach et al. (Heimbach et al., 2023b) quantified and analyzed hard-fork arbitrage in lending protocols. The tokens transferred by arbitrageurs to centralized exchanges were valued at more than \$13 million, and were actually withdrawn from the platform lenders during the Merge of Ethereum. This study complements the analysis of EIP-4844 and provides new observations on network efficiency in the interaction between builders and rollups.
Builder Behaviors.  Several previous works, such as (Öz et al., 2023; Ovezik et al., 2024; Gupta et al., 2023; Wang et al., 2025), investigated builder behaviors in PBS. The landscape exhibited that builders had a significant trend of centralization (Heimbach et al., 2023a). Likewise, Wahrstatter et al. (Wahrstätter et al., 2023) elucidated how vertical integration could potentially extract more value and further centralize block construction. Öz et al. (Öz et al., 2024) illustrated the relationship between the profit and market share, and how private order flow intensified the centralization of the builder market. In addition, Yang et al. (Yang et al., 2024) focused on competition and efficiency in MEV-Boost auctions of builders, and Heimbach et al. (Heimbach et al., 2024) quantified the non-atomic arbitrage utilized by builders to profit in the block-building process. Aside from this, blockchain censorship was defined to measure the censorship of builders in (Wahrstätter et al., 2024; Wang et al., 2023), demonstrating neutrality and security pitfalls. However, no prior works explored builder strategies from the perspective of block sizes and building choices. To the best of our knowledge, we are the first to measure builders’ strategies towards distinct transactions and evaluate their efficiency after EIP-4844, offering crucial insights for builders to refine their block construction.
Rollup Analysis.  Extensive studies on the performance of the rollup ecosystem have revealed crucial insights into its present dynamics and potential future advances (Palakkal et al., 2024; Chaliasos et al., 2024). The process of MEV extraction within Layer-2 rollups has been analyzed by Torres et al. in (Torres et al., 2024), while Avil et al. (Khalil and Rahman, 2024) investigated the arbitrage of NFT that occurs in rollups. Gogol et al. (Gogol et al., 2024b) examine the effects of inscriptions on rollups. They find that the rise in meme transactions leads to reduced gas fees for rollups. Furthermore, Stefanos et al. (Chaliasos et al., 2024) provided insight into zero-knowledge rollups and developed a structured approach for their evaluation. The research presented in (Crapis et al., 2024) introduced an economic framework to examine EIP-4844 and its impact on rollups for reducing the cost of mainnet data. Our research provide the first rollup behavior evaluation and offers our insights on efficient strategies, which is useful for rollups to optimize their strategies in the post-EIP-4844 era.
8. Conclusion
The introduction of EIP-4844 has fundamentally reshaped the economic landscape of Ethereum, creating a new marketplace for data availability. This paper offers the first comprehensive empirical and theoretical analysis of this post-upgrade environment, focusing on the evolving interactions between block builders and rollup operators. Leveraging a novel large-scale dataset encompassing over 319 million transactions, we identified key inefficiencies in current market behavior: builders frequently waste their block space for selecting less profitable type-3 transactions, and rollups often engage in redundant or suboptimal blob submission strategies. These practices result not only in significant direct financial losses, but also in broader network externalities, such as increased inclusion delays and resource congestion.
To address these challenges, we developed a formal game-theoretic framework capturing the strategic tensions at play. From this model, we derived provably efficient strategies: a new pricing rule for builders to ensure profitable block construction, and a bundling principle for rollups that minimizes redundant fees. Our findings contribute to a deeper understanding of the economic implications of Ethereum’s evolving infrastructure and offer actionable guidance to optimize outcomes for both builders and rollup operators.